A farsighted professor can read the State News at d = 175 cm, but no closer. What power (in diopters) lens should he use for reading?
Answer #1

let's assume the distance of object kept for reading is kept at distance of 25cm, as it is not given in question.

The real object is 25 cm away from the front of the lens, and we want to produce an image that appears to be 100 cm away 175cm from the lens so this will be negative. Plug everything into the equation:

power of lens (D) 1/f = 1/d1 + 1/d2

d1 = .25 m (for reading)

d2 = 1.75m

1/f = 1/.25 - 1/1.75 = 4 - .5714 = 3.4285 ~ 3.43 D